OpenStudy (anonymous):

The height of each story of an office building is 9 3/4 feet. The height is measured from the ceiling to the floor. The building has three underground levels for parking. A car is parked two levels below ground. At what elevation are the car’s tires relative to the ground level? Assume the ground level is at zero feet. A) -19 1/2 feet B) -12 feet C) 8 feet D) 19 1/2 feet

OpenStudy (anonymous):

it'd either a or b

OpenStudy (anonymous):

because it would be negative

OpenStudy (johnweldon1993):

If each level is \(\large 9\frac{3}{4}\) ft and the car is 2 levels below ground... \[\large 9\frac{3}{4} + 9\frac{3}{4} = ?\]